A proton moves through a uniform magnetic field of magnitude 2.0 μT at a speed of 5.00x10^6 m/s.


(a) What is the magnitude of the maximum magnetic force it can experience?

(b) If the proton is replaced by an electron, will it experience the same maximum force?


1
Expert's answer
2022-05-02T13:43:04-0400

(a)

"F=qvB\\
=1.6*10^{-19}*5.00*10^6*2.0*10^{-6}\\
=1.6*10^{-18}\rm \: N"

(b) the magnitude of the electron's charge is the same as proton, so the force would be the same.